Session: Sep 2000 (session 3)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (A.Scalambra at Tue Sep 26 22:25:31 METDST 2000) Success

Shanghai (Huang Xinyong at Wed Sep 27 03:51:17 METDST 2000) Success
No any problems.

Torun (K.Borkowski at Wed Sep 27 09:54:04 METDST 2000) Minor failures
FS reported "no_1pps" in BBC7 and (a few times only) in BBC6. Otherwise no known problems.

Noto (C.Trigilio at Fri Sep 29 15:40:11 METDST 2000) Failure
Did not observe because of a failure of the tape recorder (capstan motor) occurred on 22 Sep.

Onsala (John Conway at Fri Sep 29 15:51:04 METDST 2000) Success
no known problems

Effelsberg (A. Kraus at Wed Oct 4 15:43:11 METDST 2000) Minor failures
The observations had to be interrupted between UT1255-1415 due to technical problems. Until UT0725 we used ifd=nor,nor instead of ifd=nor,alt.

Jodrell_Bank (Alastair Gunn at Thu Oct 5 09:59:57 METDST 2000) Minor failures
First scan missed due to tape loading problems. Extreme RFI as usual.

Cambridge (Alastair Gunn at Thu Oct 5 10:00:27 METDST 2000) Minor failures
FS crashed occasionally and was stopped to check bandswitch control problems. As far as we know the data should not have been affected.

\(Other\) ( at Sat Jul 7 13:04:28 METDST 2001) Minor failures